OpenStudy (anonymous):

factor: x^3-5x^2-9x+45

OpenStudy (cwrw238):

factor by grouping: : x^3-5x^2-9x+45 = x^2(x - 5) -9(x - 5) = (x - 5)(x^2 - 9) can you continue by factoring x^2 - 9?